40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Apr 5 07:19:16 2012, kiwamu, Update, LSC, DRMI locking DRMI_2012Apr4_edit.png
    Reply  Sat Apr 7 01:56:05 2012, kiwamu, Update, LSC, OSA signal in DRMI condition OSA.png
       Reply  Sat Apr 7 06:58:34 2012, kiwamu, Update, LSC, DRMI lock : lost good alignment 
Message ID: 6506     Entry time: Sat Apr 7 01:56:05 2012     In reply to: 6489     Reply to this: 6508
Author: kiwamu 
Type: Update 
Category: LSC 
Subject: OSA signal in DRMI condition 

It wasn't a dream or illusion -- I was locking the DRMI to the right condition last Wednesday (#6489).

Here is a snap shot of the AS-OSA signal taken today when the DRMI was locked with the same control settings (#6489).

The blue curve is data taken when the PRMI was locked for comparison.

You can see that both the upper and lower 55 MHz sideband are amplified by the SRC.

OSA.png

 

(Some notes)

Currently SRM is slightly misaligned such that the MICH optical gain at AS55Q doesn't increase so much with the presence of SRM.

With this condition I was able to acquire the lock more frequently than how it used to be on the Wednesday.

The next step is to gradually align SRM, to optimize the controls and to repeat this process several times until SRM is fully aligned.

Quote from #6489

A first thing I tried was : flipping the control sign of the SRCL while keeping the same control setups for the PRCL and MICH.

Occasionally the DRMI was "sort of" locked and hence I believe this setup must be a good starting point.

ELOG V3.1.3-